OpenStudy (anonymous):

find the limit of sin^2x/4x as x approaches pi/6

OpenStudy (kainui):

Just plug it in.

OpenStudy (agreene):

\[\large \lim_{x \rightarrow \frac{\pi}6}\frac{\sin^2 x}{4x}\] if you notice, the only place for this function to be undefined is when the denominator is 0; the only case for this is x=0; other than that... you can just plug it in

OpenStudy (solomonzelman):

Factor out of 1/4. then take the limit of the numerator and denominator separately.

OpenStudy (solomonzelman):

I would prefer that you post your work in here.